OpenStudy (anonymous):

calculate the ratio of the second term to the first term, and express the answer as a decimal value. State the percent increase or decrease a. 80,60 b. 40,45 c. 88,198 D, 36,32 e. 110,96.8 f. 63,100.8

OpenStudy (anonymous):

60:80 60/80=0.75, 60=75% of 80 60:80 = 0.25:0.75 60+x=80 x=20 20/60 = (100/3)%, 20 = (100/3)% of 60 (100%) + (100/3)%= (400/3)% therefore 60:80= 0.25:0.75, the change is + (400/3)%